What is the force constant for the suspension system of a car that settles 1.5 cm when an 85 kg person gets in?
Answer #1

here,

mass of person , m = 85 kg

the compression , x = 1.5 cm = 0.015 m

let the force constant be K

equating the forces

K * x = m * g

K * 0.015 = 85 * 9.81

solving for K

K = 5.56 * 10^4 N/m

the force constant of the suspension system is 5.56 * 10^4 N/m